OpenStudy (anonymous):

(Please help) Let f(x) = 3x2 + x - 3 and g(x) = x2 - 5x + 1. Find f(x) - g(x). 2x2 - 4x - 2 2x2 - 4x - 4 2x2 + 6x - 2 2x2 + 6x - 4

OpenStudy (anonymous):

f(x) - g(x) = (3x^2 + x - 3) - (x^2 - 5x +1) = (3x^2 + x - 3) - x^2 + 5x - 1 = 2x^2 + 6x - 4.

OpenStudy (anonymous):

Thank you so much! Do you think you could help with just one more?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Subtract (2x3 + 9x - 8) - (4x2 - 15x + 7). 2x2 - 6x - 1 -2x2 + 24x - 15 2x3 - 4x2 - 6x - 1 2x3 - 4x2 + 24x - 15

OpenStudy (anonymous):

Do you think you could explain how it's done a bit?

OpenStudy (anonymous):

First distribute the negative between the two equations into the second equation. So it becomes (2x^3 + 9x - 8) - 4x^2 + 15x - 7. Next, subtract the similar variables! -2x^2 + 24x - 15.

OpenStudy (anonymous):

Ok I understand and sort of get it, thank you for the help :)

OpenStudy (anonymous):

No problem! :)